OpenStudy (anonymous):

solve inequality solve interval notation: 5x+4>4 and 3x-7<-1 anyone plzz

hero (hero):

The All-Pleasing Eye?

OpenStudy (anonymous):

lol its my eye!

hero (hero):

lol

OpenStudy (anonymous):

5x+4>4 => 5x > 0 => x > 0 3x-7<-1 3x<6 x<2 0<x<2

OpenStudy (anonymous):

\[5x+4\ge4 and 3x-7\le-1\]

OpenStudy (anonymous):

i cudnt properly input symbols out there..

OpenStudy (waheguru):

http://openstudy.com/#/updates/4edc2527e4b01b4ec492850d hi GT

OpenStudy (anonymous):

0 <= x <= 2

OpenStudy (anonymous):

ur right GT!! thnx

OpenStudy (anonymous):

\[8x+5\ge5 and 3x-7\ge-1\]

OpenStudy (anonymous):

im tryna follow your lead..

OpenStudy (anonymous):

is it [2, 2]

OpenStudy (anonymous):

or 0, 2

OpenStudy (anonymous):

yes 0, 2

OpenStudy (anonymous):

[0,2]

OpenStudy (anonymous):

\[2s+3<13 or -2s+0\le-7\]

OpenStudy (anonymous):

not 0 sorry its a 9

OpenStudy (anonymous):

is it 5, 0..?

OpenStudy (anonymous):

nope nt 5, 0